OpenStudy (marihelenh):

The gold foil used by Ernst Rutherford in his investigations 1 um thick. Given that the radius of gold is approximately 130 pm, how many atoms thick was the foil?

OpenStudy (aaronq):

How many atoms fit across the foil (assuming they're liked up side by side) \(1~\mu m=1.0*10^{6}~pm\)

OpenStudy (marihelenh):

\[\frac{ 130p m }{ 1 } * \frac{ 1\mu m }{ 1.0*10^{6}p m }\]

OpenStudy (marihelenh):

Is this how I would set that part up?

OpenStudy (marihelenh):

@aaronq

OpenStudy (aaronq):

yes that would give you the size of one atom in micrometers, then you divide 1 micrometer (the thickness of the foil) by the size of 1 atom (in micrometers)